OpenStudy (anonymous):

reviewing for exam...can someone help me work through this? Solve.... x-y+z=0 3x-2y+6z=9 -x+y-2z=-2

OpenStudy (anonymous):

x - y + z = 0 -->(-3) 3x - 2y + 6z = 9 ----------------- -3x + 3y - 3z = 0 (result of multiplying by -3) 3x - 2y + 6z = 9 -----------------add y + 3z = 9 3x - 2y + 6z = 9 -x + y - 2z = -2 -->(3) --------------- 3x - 2y + 6z = 9 -3x + 3y - 6z = -6 (result of multiplying by 3) --------------add y = 3 y + 3z = 9 3 + 3z = 9 3z = 9 - 3 3z = 6 z = 2 x - y + z = 0 x - 3 + 2 = 0 x - 1 = 0 x = 1 check... 3x - 2y + 6z = 9 3(1) - 2(3) + 6(2) = 9 3 - 6 + 12 = 9 -3 + 12 = 9 9 = 9 (correct) x = 1, y = 3 and z = 2

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

anytime :)